[Brahms-l] Shift report 20040215 08:00-16:00

From: Apache <apache@bnl.gov>
Date: Sun Feb 15 2004 - 15:53:36 EST
------------------------------------------------------------
Date: Sun Feb 15 15:53:36 2004, Shift: 08:00-16:00
Supervisor: DO, with: 
Log book#15 pages 58 - 58

Shift Summary:
------------------------------------------------------------
1.Almost no problem, they lost steering during run 10373, recovered quickly. Had two nice stores today: beam hours = 5.5 hours

2.Data: MRS run plan is achieved. Flemming decided to run
  the MRS at 45deg (cannot move remotely) at A1050.
  
  Runs        MRS tr3        FS tr2 (4deg A 1/1)
  10369        138k              63k 
  10374        183k             150k 
  10375        130k              85k
  10376        117k              71k
  10377        still running

Pedestals: non TPC 10370, TPC 10371
 
3.Well, nothing to declare :) Oh yes, TRFS top TDC 7 is back...
------------------------------------------------------------

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Sun Feb 15 15:53:53 2004

This archive was generated by hypermail 2.1.8 : Sun Feb 15 2004 - 15:54:10 EST